OpenStudy (anonymous):

Can you guide me to prove the following theorem? Thanks in advance! Secant-Tangent Intersection Theorem What it says: When a secant and tangent intersect at the point of tangency, the angles created at the point of intersection are half the measurement of the arcs they intersect. What it means: Imagine a line crosses through a circle (secant) and another line runs alongside the circle (tangent). These lines meet at a point on the circle. The two angles at this point are half the measure of the arcs they create on the circle.

OpenStudy (anonymous):

What it looks like: http://i.imgur.com/IgfcqpT.png

ganeshie8 (ganeshie8):

Join \(SN\), join \(NL\)

ganeshie8 (ganeshie8):

Look at triangle, \(\triangle SNL\)

ganeshie8 (ganeshie8):

can u find the angle, \(\angle SNL\) in terms of the central angle, \(\angle NSL\) ?

ganeshie8 (ganeshie8):

u may use "triangle sum property" and "isosceles triangle"

ganeshie8 (ganeshie8):

Since \(\triangle SNL\) is an isosceles triangle, its base angles are cnongruent : \(\large \angle SNL \cong \angle SLN\)

OpenStudy (anonymous):

Okay. Since all of the angles in a triangle add up to 180, and since angle SLN has to be congruent to angle SNL because it is isosceles, then it would be (180-(angle NSL))/2.

ganeshie8 (ganeshie8):

By triangle sum property we have : \(\large \angle SNL + \angle SLN + \angle NSL = 180\)

ganeshie8 (ganeshie8):

Yes ! you have it, \(\large 2\angle SNL = 180 - \angle NSL\) \(\large \angle SNL = \frac{180 - \angle NSL}{2}\)

ganeshie8 (ganeshie8):

fine, so far eh ?

OpenStudy (anonymous):

yeah, i think i have an idea of where you're going to take this

OpenStudy (anonymous):

but im not completely sure, so lets go on

ganeshie8 (ganeshie8):

good :)

ganeshie8 (ganeshie8):

Next, we knw that tangent and radius intersect at 90 degrees, so we have \(\large \angle SNE \cong 90\)

ganeshie8 (ganeshie8):

however \(\large \angle SNE \cong \angle SNL + \angle LNE\)

ganeshie8 (ganeshie8):

so, \(\large 90 \cong \frac{180 - \angle NSL}{2} + \angle LNE \)

ganeshie8 (ganeshie8):

you can solve \(\angle LNE\)

OpenStudy (anonymous):

Okay. \[180=(180-\angle NSL)+2\angle LNE \\ 0=-\angle NSL + 2\angle LNE \\ \angle NSL = 2\angle LNE \\ \angle LNE = \frac{ 1 }{ 2 }\angle NSL\]

OpenStudy (anonymous):

wow, thank you

ganeshie8 (ganeshie8):

looks great ! u wlc :)